# SlowPoke: Understanding and Detecting On-Chip Fail-Slow Failures in Many-Core Systems

### v1

@@ 4.3. Data Collection Overhead
In this experiment, we evaluate the runtime overhead introduced by the proposed probe-based data collection mechanism. We simulate the execution of DarkNet-19, ResNet-50, GoogLeNet, VGG and Binary Tree workloads under different probe configurations. In the simulator, the latency of a probe is determined by the amount of data it collects. For example, retrieving the current system clock incurs a 10-cycles delay. We use the total execution time without any probes as the baseline. The Communication Probe, Computation Probe, and Full Probe configurations correspond to inserting probes into communication instructions, computation instructions, and all instructions, respectively, in order to assess the overhead in different analysis scenarios. Considering the variability in hardware performance, we repeat the experiments multiple times under the same configuration and report the average execution time as the measured overhead.
Figure 10. Time overhead of SL-Compiler probing.
Figure 10 shows that our probe-based data collection mechanism gathers runtime traces with low overhead. Even in the most intrusive configuration, where probes are inserted into all instructions, the execution time increases by at most 10% compared to the baseline. For more targeted scenarios, (e.g., Communication Probe for network performance monitoring), the overhead remains below 5%.
4.4. Effectiveness of Storage Overhead Reduction
In this experiment, we evaluate the storage effectiveness of SL-Recorder in reducing the storage overhead of trace data through compression. We use the five workloads mentioned before to measure the memory consumption of trace data under both raw and compressed formats. The raw data records most of the information such as index, timestamps and operands of an instruction, and is considered as the baseline. For each workload, we separately record the storage overhead of communication-related traces and computation-related traces, and then apply SL-Recorder’s compression mechanism to obtain the reduced storage cost. We compare the total memory consumption before and after compression to quantify the storage saving achieved by SL-Recorder.
@@ 4.4. Effectiveness of Storage Overhead Reduction
Figure 11. Memory cost of communication and computation traces with and without SL-Recorder.
Figure 11 shows the memory cost of raw and compressed traces across different workloads. In this figure, Comm represent the memory required for communication and Comp represent that of computation traces, respectively. We observe that SL-Recorder achieves substantial storage reduction and is effective across workloads of different scales. For small-scale workloads such as DarkNet-19, SL-Recorder reduces the storage overhead to 0.8% and 0.7% of the original sizes for communication and computation traces, respectively. Even for large-scale workloads such as VGG, SL-Recorder maintains a similar compression ratio, reducing a total of approximately $17.8\text{\,}\mathrm{MiB}$ of trace data to only $137.2\text{\,}\mathrm{KiB}$. Overall, these results demonstrate that SL-Recorder can reduce storage overhead by more than two orders of magnitude, significantly lowering memory requirements.
@@
Figure 12. SL-Recorder compression rate: Impact of the Hash (H), Bucket (B), Size (S), and Threshold (T).
@@
4.5. Design Space Exploration
To find out the optimal configuration of SL-Recorder, we conduct a design space exploration (DSE) across four key parameters: the number of hash functions (Hash), the number of buckets (Bucket), the Stage-2 size (Size), and the pattern-selection threshold (Threshold). These parameters jointly determine the compression efficiency and detection accuracy of SL-Recorder.
@@ 4.5. Design Space Exploration
Figure 13. SL-Recorder Design Space Exploration.
Figure 13(a) shows the exploration results on a $\times4$ many-core architecture, where each point corresponds to a parameter configuration, and different colors represent different workloads. We can observe that the design space exhibits clear trade-offs: configurations with lower memory cost often suffer from degraded accuracy, whereas moderate-cost configurations yield balanced performance. The lowest-cost Pareto-optimal cases are highlighted with red triangles. Figure 13(b) shows DSE results on a $\times8$ many-core architecture, which has the same trade-off like $\times4$ architecture.
4.6. Scalability Analysis
To evaluate the scalability of SlowPoke, we test its performance on many-core architectures of increasing scale. Specifically, we simulate ResNet-50 and DarkNet-19 on $\times4$, $\times6$, and $\times8$ architectures, and measure the execution time under three probing modes: communication probe, computation probe, and full probe.
The results are shown in Figure 14. For both networks, the overall execution time decreases as the architecture scale grows, which indicates that larger many-core systems provide higher parallelism and better workload distribution. More importantly, the additional overhead introduced by probing remains consistently low across different scales. As illustrated by the green points in the figure, the normalized execution time overhead of the full probe is always below 10%, demonstrating that our framework introduces only marginal overhead even at larger scales.
Figure 14. Scalability of SlowPoke: Time overhead.
Figure 15. Scalability of SlowPoke: Memory cost and detection accuracy.
@@ 4.6. Scalability Analysis
5. Related Work
Distributed tracing and storage. Prior studies reduce tracing overhead through selective collection and compact storage. Mint (Huang et al., 2025) and Pivot Tracing (Mace et al., 2015) lower cost via commonality analysis and dynamic causal monitoring, while Hindsight (Zhang et al., 2023) and Sieve (Huang et al., 2021) apply selective tracing and attention-based sampling. Lightweight sketches such as Dynamic Sketch (Wang et al., 2019b) and Stable-Sketch (Li and Patras, 2024) offer compact representations for stream data. However, these methods generally preserve high-frequency items statically once stored. In contrast, our SL-Recorder adaptively updates patterns and retains traces most relevant to fail-slow behaviors.
Performance failure detection. Many works have analyzed hardware-induced performance variance in large-scale systems. ScalAna (Jin et al., 2020) and What-if Analysis (Lin et al., 2025) employ graph analysis for bottlenecks, while vSensor (Tang et al., 2018) and Vapro (Zheng et al., 2022b) use workload snippets to detect fluctuations. Profiling-based tools like Scalasca (Wolf et al., 2008) and GVARP (You et al., 2024) also target variance in parallel applications. Recent systems, e.g., FALCON (Wu et al., 2024), isolate stragglers in cloud and training workloads. However, these methods assume abundant storage, which is not suitable for on-chip environments. Our SL-Tracer works under strict memory limits while pinpointing fail-slow components.
Root cause analysis. Root cause analysis (RCA) often leverages structured traces and graph models for localization. Event-graph approaches like Groot (Wang et al., 2022) and GNN-based Sleuth (Gan et al., 2024) capture causal anomalies, while Sage (Gan et al., 2021) and MicroRank (Yu et al., 2021) apply spectrum analysis and performance metrics. Systems such as IASO (Panda et al., 2019) and TraceRCA (Li et al., 2021) detect fail-slow symptoms in distributed services. However, these works focus on communication dependencies, overlooking hardware topology. Our SL-Tracer unifies both communication and hardware structures for more accurate fail-slow root cause localization.
6. Conclusion

### v2

@@ IV-C Data Collection Time Overhead
In this experiment, we evaluate the runtime overhead of probe-based data collection under different instrumentation configurations. In the simulator, probe latency is determined by collected data volume—for example, reading the system clock incurs a 10-cycle delay. We use probe-free execution as the baseline and compare three configurations: Communication Probe (instruments communication instructions only), Computation Probe (instruments computation instructions only), and Full Probe (instruments all instructions). Each configuration is repeated multiple times to account for performance variability, with average execution time reported.
Figure 10 shows that probe-based collection introduces minimal overhead. Even under the most extensive instrumentation (Full Probe), execution time increases by at most 10% across all workloads. For targeted scenarios such as Communication Probe—suitable for network performance monitoring—overhead remains below 5%. These results demonstrate that our approach enables fine-grained trace collection with negligible performance impact.
IV-D Effectiveness of Storage Overhead Reduction
We evaluate SL-Recorder’s storage efficiency by comparing memory consumption before and after compression. The raw format records full instruction details (index, timestamps, operands, …) and serves as the baseline. For each workload, we separately measure communication and computation trace sizes, then apply compression to quantify storage savings. We also compare against IASO (communication traces), Perseus and ADR (computation traces)’ memory requirement to demonstrate SL-Recorder’s advantage.
@@ IV-D Effectiveness of Storage Overhead Reduction
Figure 11: Memory cost of communication traces with no optimization, IASO and SL-Recorder.
Figure 12: Memory cost of computation traces with no optimization, Perseus, ADR and SL-Recorder.
Figure 11 shows that SL-Recorder achieves significant compression compared to both raw traces and IASO. For Binary Tree, raw traces consume $10\text{\,}\mathrm{MiB}$ and IASO consume $7\text{\,}\mathrm{MiB}$, while SL-Recorder compresses to $86\text{\,}\mathrm{KiB}$—a 98.8% reduction. Similar results appear across all workloads. IASO’s algorithm only provides marginal savings (30-40% reduction) as it still retains full communication trace. Figure 12 demonstrates similar effectiveness for computation traces. Across all workloads, SL-Recorder consistently achieves 98-99% compression ratios, while Perseus and ADR achieve only 25-50% compression as they retain per-instruction records for regression analysis or adaptive threshold. Overall, SL-Recorder reduces both communication and computation trace storage by over two orders of magnitude across all workloads, significantly outperforming existing approaches and making fine-grained trace collection practical for resource-constrained environments.
@@
Figure 13: SL-Recorder compression rate: Impact of the Hash (H), Bucket (B), Size (S), and Threshold (T).
@@
Figure 14: SL-Recorder Design Space Exploration.
IV-E Design Space Exploration
To find out the optimal configuration of SL-Recorder, we conduct a design space exploration (DSE) across four key parameters: the number of hash functions (Hash), the number of buckets (Bucket), the Stage-2 size (Size), and the pattern-selection threshold (Threshold). These parameters jointly determine the compression efficiency and detection accuracy of SLOTH.
Figure 14(a) shows the exploration results on a $\times4$ many-core DNN accelerator, where each point corresponds to a parameter configuration, and different colors represent different workloads. We can observe that the design space exhibits clear trade-offs: configurations with lower memory cost often suffer from degraded accuracy, whereas moderate-cost configurations yield balanced performance. The lowest-cost Pareto-optimal cases are highlighted with red triangles. Figure 14(b) shows DSE results on a $\times8$ many-core DNN accelerator, which has the same trade-off like that on $\times4$ accelerator.
Figure 15: Convergence analysis of FailRank on different fail-slow cases. The y-axis (log scale) represents the L1-norm difference between score vectors of consecutive iterations. The algorithm converges rapidly within 20 iterations for all datasets ($\epsilon=10^{-4}$).
Figure 16: Scalability of SLOTH: Time overhead.
Figure 17: Scalability of SLOTH: Memory cost and detection accuracy.
IV-F FailRank Convergence Analysis
@@
Figure 15 illustrates the reduction of $\Delta_{k}$ over iterations on a logarithmic scale, revealing a distinct and rapid convergence trend. The process typically begins with an initial rapid descent phase during the first few iterations (iterations 1-3), where the error drops sharply by an order of magnitude as the ranking scores propagate through the immediate neighborhood of the failure. This is immediately followed by a stable decay phase, where the error decreases linearly on the log-scale plot, indicating a geometric rate of convergence as the global ranking stabilizes.
Crucially, the curves representing different fail-slow locations and intensities exhibit highly consistent trajectories, demonstrating that FailRank is robust to variations in failure patterns. In all tested cases, the algorithm satisfies the convergence criterion within approximately 17 iterations. Given that each iteration involves only lightweight matrix-vector operations, this low and bounded iteration count ensures that FailRank incurs negligible runtime overhead, making it well-suited for detection within the resource-constrained environment of many-core accelerators.
IV-G Scalability Analysis
To evaluate the scalability of SLOTH, we test its performance on many-core DNN accelerators of different scales. Specifically, we simulate ResNet-50 and DarkNet-19 on $\times4$, $\times6$, and $\times8$ accelerators, and measure the execution time under three probing modes: communication probe, computation probe, and full probe.
The results are shown in Figure 16. For both networks, the overall execution time decreases as the architecture scale grows, which indicates that larger many-core DNN accelerators could provide higher parallelism and better workload distribution. More importantly, the additional overhead introduced by probing remains consistently low across different scales. As illustrated by the red points in the figure, the normalized execution time overhead of the full probe is always below 10%, demonstrating that our framework introduces only marginal overhead even at larger scales.
@@ IV-G Scalability Analysis
V Related Work
Distributed Tracing. Existing distributed tracing systems reduce overhead through reactive collection or selective storage[34, 38, 37, 62], but these approaches are often not suitable for diagnosing fail-slow behaviors. For instance, Pivot Tracing [48] enables on-demand instrumentation based on user queries, yet critical early-stage evidence of gradual degradation is often missed. Hindsight [81] uses post-hoc triggers to commit buffered traces, but the performance decay of fail-slow issues may not activate these triggers in time, leading to the loss of crucial data. Similarly, while Sieve [31] applies an attention model to prioritize traces, it is biased towards conspicuous spikes and can overlook the weak, temporally-correlated signals characteristic of a fail-slow. In contrast, our SL-Compiler employs a proactive and structure-aware strategy; by analyzing the DNN’s computational graph to automatically instrument critical dataflows (while also allowing user guidance), it can capture most of the traces indicating fail-slow degradation.
Trace Storage Optimization. The sheer volume of trace data needs optimizations for efficient storage[43, 30, 58]. For example, Mint [30] reduces storage footprint by decomposing traces into common patterns and specific attributes. However, this structural compression is agnostic to performance semantics; it does not help in identifying or prioritizing traces that exhibit gradual latency degradation. Other works leverage sketches for data summarization[39, 75, 51]. For example, Dynamic Sketch [75] uses a doorkeeper mechanism to efficiently detect heavy hitters, but it focus on request frequency and is misaligned with fail-slow analysis. Similarly, while SketchLib [53] provides highly optimized sketch implementations for Reconfigurable Match-Action Tables (RMT), its design is incompatible with the stringent on-chip memory constraints. Our SL-Recorder employs a novel two-stage sketch design. By extending traditional sketch features to include latency statistics, it can identify similar workloads within a DNN and dynamically update them to filter the traces that are most indicative of fail-slow behavior.
Failure Detection and Root Cause Analysis. Ensuring system reliability requires robust mechanisms for both failure detection and subsequent root cause analysis (RCA)[71, 21, 20, 33]. For instance, ScalAna [33] leverages backtracking on a Program Structure Graph (PSG) to localize root causes. However, its reliance on a purely logical graph, which abstracts away the hardware topology, hinders the diagnosis of communication-related bottlenecks. Another class of methods focuses on component-level metrics[55, 46, 45]. Approaches like IASO[55], using timeout signals and peer-scoring, and ADR [45], which monitors latency and frequency shifts, are effective at identifying node-level slowdowns. Nevertheless, their node-centric perspective fails to capture performance degradation occurring on the communication links themselves. Our SL-Tracer overcomes these limitations using a Multi-level Communication Graph (MCG) that integrates software communication dependencies with the underlying physical topology. By applying an iterative FailRank algorithm on MCG, SL-Tracer can pinpoint the root cause of fail-slow, irrespective of whether the origin is a computational node or a communication link.
VI Conclusion
[…]